# Estimating velocities of infectious disease spread through spatio-temporal log-Gaussian Cox point processes

### v1

@@ Abstract
Understanding the spread of infectious diseases such as COVID-19 is crucial for informed decision-making and resource allocation. A critical component of disease behavior is the velocity with which disease spreads, defined as the rate of change between time and space. In this paper, we propose a spatio-temporal modeling approach to determine the velocities of infectious disease spread. Our approach assumes that the locations and times of people infected can be considered as a spatio-temporal point pattern that arises as a realization of a spatio-temporal log-Gaussian Cox process. The intensity of this process is estimated using fast Bayesian inference by employing the integrated nested Laplace approximation (INLA) and the Stochastic Partial Differential Equations (SPDE) approaches. The velocity is then calculated using finite differences that approximate the derivatives of the intensity function. Finally, the directions and magnitudes of the velocities can be mapped at specific times to examine better the spread of the disease throughout the region. We demonstrate our method by analyzing COVID-19 spread in Cali, Colombia, during the 2020-2021 pandemic.
Keywords: Bayesian inference, Stochastic Partial Differential Equations, Finite differences
@@ 1 Introduction
Transmission of infectious diseases such as COVID-19, as we saw in recent years, has the potential to collapse the healthcare system, causing health, societal, and economic problems [Kaye et al., 2020] [Wu et al., 2021]. For that reason, modeling these diseases is critical to understanding how they spread, and this could help decision-makers create policies to mitigate them. In this regard, a powerful tool in the context of spatio-temporal point patterns that can help us understand the spread of an infectious disease is the velocity that could be computed given a direction in any location in space and time. This magnitude could be understood as the instantaneous rate of change for the point pattern at location and time $(\textbf{u},t)$ in direction v.
In the context of point patterns, velocity is a quantitative measure that helps us to understand how a spatio-temporal process evolves. Moreover, when it is possible to define the intensity function, velocity refers to how fast the intensity changes in time relative to its change in a given spatial direction. Thus, this measure captures the speed of the intensity change in that direction and provides information about the spatio-temporal dynamics of intensity and, consequently, the underlying point pattern. In this context, velocity could be interpreted in an instantaneus sense as the chance for an event to lie in a small product neighborhood of a specific location and time.
@@
In this paper, we present how to compute the minimal velocity for infectious diseases. However, a different potential application is using point patterns of species location to understand the direction and velocity of their movements [Moraga, 2020]. That could help understand their distribution in some specific area related to climate change or other phenomena. Another application is crime data in a given city, which could help us understand how crime changes over time.
@@
Many authors implemented different models to estimate the directional derivatives of the intensity of a point pattern. For instance, [Banerjee and Gelfand, 2003] introduced a formal directional derivative process based on the realization of a stochastic process. Later, [Terres and Gelfand, 2016] developed a methodology for local sensitivity analysis based on directional derivatives that also uses continuous covariates to relate them with spatial gradients. For geostatistical data, temporal and spatial derivatives were computed to estimate velocities of climate change [Schliep et al., 2015] relating covariates to this, for example, temperature. Extending these ideas to point patterns [Schliep and Gelfand, 2019] compute velocities for spatio-temporal point patterns, assuming that the point pattern follows a log-Gaussian Cox (LGCP) process and also by using Nearest Neighbor Gaussian process (NNGP). Additionally, other authors consider other ways to compute velocities that use higher-order derivatives, such as second-order derivatives [Quick et al., 2015].
The main contribution of this paper is computing the minimal velocity using finite differences to approximate partial derivatives of the intensity function. The principal advantage is that we are not using a parametric form of the intensity function or a closed form of the covariance structure. This gives more flexibility to the intensity function, which is an LGCP as in [Schliep and Gelfand, 2019], but we are not fixing the smoothing parameter. This procedure also allows us to capture instantaneous changes that parametric models are slow to capture over time. For the computational part, we use the integrated nested Laplace approximation (INLA) [Rue et al., 2009] that provides computational efficiency to compute the parameters of the parameters of the intensity function.
@@
The structure of the remainder of the paper is as follows. In Section 2, we introduce some theory about point processes and the log-Gaussian Cox process. Then, in Section 3, we fully detail how we decompose the intensity function into three components and how the inference for the LGPC is defined by its covariance structure. In this section, we also explain how velocity is defined for spatio-temporal point patterns and how to approximate it using finite differences. In Section 4, we present a simulation study to analyze the performance of our method using a point pattern with a deterministic intensity function. Then, in Section 5, we present an application using data from COVID-19 in Cali, Colombia. Finally, in Section 6, we present some conclusions, a discussion of the findings, and some directions for future research.
@@ 2 Preliminary concepts
In this section, we first present some theoretical concepts related to point processes necessary to introduce the methodology used to model the spread of disease, and to compute velocities. Using these concepts, we can provide a detailed explanation of how the spread of disease can be modeled both spatially and temporally. This involves analyzing the locations of infected individuals as they change over time. By tracking these observed locations over a specific period, we can gain valuable insights into the patterns and dynamics of disease transmission. This approach allows us to understand where the infections occur and how they evolve, which is crucial for effective disease control and prevention.
@@ 2.1 Point processes
Given a subset $W\subseteq\mathbb{R}^{2}$ and an interval $T\subseteq\mathbb{R}$, a spatio-temporal point pattern consists of a collection of all points $\{\zeta_{i}\}_{i=1}^{n}=\{\left(\textbf{u}_{i},t_{i}\right)\}_{i=1}^{n}\subseteq W\times T$. Under these conditions, a spatio-temporal point process is a random mechanism whose outcome or realization is a point pattern. For a spatio-temporal point process we define the intensity function $\lambda:W\times T\to[0,\infty)$ such that $\int_{A}\int_{B}\lambda(\textbf{u},t)\mathrm{d}t\mathrm{~{}d}\textbf{u}<\infty$ for all bounded sets $A\subseteq\mathbb{U}$, and intervals $B\subseteq\mathbb{T}$ that satisfies the following property
@@
One of the most important point processes is the Poisson process. This process is frequently used due to its simplicity, and satisfies the following conditions.
@@ 2.2 Log-Gaussian Cox processes
Log-Gaussian Cox processes (LGCPs) are typically used to model complex relationships between space and time, such as the spread of a contiguous disease [Diggle et al., 2013]. To understand this complex process, we start by discussing the Cox process. A Cox process can be viewed as a Poisson process with a random intensity function. That is, the intensity function $\Lambda(\textbf{u},t)$ meets these two postulates
@@
    $\Lambda(\textbf{u},t)$ with $\textbf{u}\in W$ and $t\in T$ is a non negative valued stochastic process,
@@
LGCP models were introduced as a particular case of a Cox process [Møller et al., 1998]. As the name suggests, an LGCP is a Cox process with intensity function given by a Gaussian process, that is $\log(\Lambda(\mathbf{u},t))=\mu(\textbf{u},t)+\xi(\mathbf{u},t)$, where $\exp(\mu(\textbf{u},t))$ could be interpreted as the mean structure of $\Lambda(\textbf{u},t)$ and $\xi(\mathbf{u},t)$ is a stationary Gaussian process.
@@ 3 Methodology
To estimate velocities of an infectious disease, we propose a spatio-temporal point process model to describe how infectious diseases spread in space across time.Specifically, we use a LGCP model where the intensity is multiplicatively decomposed into temporal, spatial, and spatio-temporal stochastic variation as presented in [Diggle et al., 2005] and [Ribeiro Amaral et al., 2023]. Once the model is fitted, the intensity estimate is used to compute minimal velocities through finite differences. This approach allows us to understand the direction and velocity of disease spread.
@@ 3.1 Spatio-temporal LGCP model
$\Lambda(\textbf{u},t)=\eta(\textbf{u})\cdot\mu(t)\cdot\xi(\textbf{u},t).$ (1)
Here, the spatial and temporal components are deterministic and capture aspects of the underlying population at risk of the pattern of disease. The spatial component $\eta(\textbf{u})$ represents the population density in the study region. This implies that if there were zones or locations with higher or lower cases, this variation would be absorbed in $\eta(\textbf{u})$. The temporal component $\mu(t)$ is estimated by grouping the total number of cases over time, and using a spline model with different basis functions to model this data. Thus, the flexibility of the splines estimation of $\mu(t)$ could capture the anomaly in the number of cases across time. This emphasizes that the remaining term $\xi(\textbf{u},t)$ will capture anomalies in space and time. This term is modeled as a stationary, zero-mean log-Gaussian stochastic process. In the following sections, we explain in detail how we compute each component and their importance in the model.
@@ 3.1.1 Temporal component
The first step in our approach is to model the number of individuals per temporal unit (e.g., days, hours, etc.) infected by the disease. Accurately capturing this data is crucial for understanding the spread and making reliable predictions. We employed a spline model, a flexible statistical tool capable of fitting complex data patterns to achieve this goal.
Splines allow us to handle non-linear trends in the data, providing a more precise fit than traditional linear models [Wood, 2017]. This flexibility is essential in epidemiological studies, where the number of infections can vary greatly due to numerous factors such as periodic changes due to day-of-week effects or time-of-year events. We, therefore, estimate the number of infected people in time $t$, denoted by $\mu(t)$ by fitting the following standard log-spline model
@@
where $b_{i}$ are completely known functions, commonly referred to as basis functions and $q$ denotes the number of functions. Some examples of these basic functions are polynomials, roots, cosine, sine, and step functions. However, the election of the basic functions and their number highly depends on the data.
@@ 3.1.2 Spatial component
Different methods could be used to compute the population density $\eta(\textbf{u})$ over a study region. In this work, we use population information from WorldPop [Bondarenko et al., 2020], an open-access project that combines official census information and satellite images to estimate population density at different resolutions. For regions where population information is not available, we could estimate the population density using nonparametric methods such as Gaussian kernel estimators [González and Moraga, 2023]. For instance, the population density could be estimated as
@@
where $h>0$ is a fixed bandwidth, $\phi(\textbf{u})=(2\pi)^{-1}\exp{\{-0.5\lVert\textbf{u}\rVert^{2}\}}$, and $u_{i}$ are the location of the infected individuals in the study region across time. We could also use an adaptive bandwidth $h_{i}$ for each location $\textbf{u}_{i}$ to have a better approximation.
@@ 3.1.3 Spatio-temporal component
Once the spatial and temporal components are estimated, they can be used to compute the intensity function of the spatio-temporal process. In this case, we assume that the observed point process is generated by a log-Gaussian Cox process. Then, the counting number of infected individuals in the region $W$ at the interval time $T$ called $\mathcal{N}\left(W,T\right)$ meets the following conditions
@@
and the intensity function described in Equation (3) could be transformed as
$\Lambda(\textbf{u},t)=\eta(\textbf{u})\cdot\mu(t)\cdot\exp\left(\zeta(\textbf{u},t)\right).$ (3)
Here, $\zeta(\textbf{u},t)$ is a stationary Gaussian process with mean $\mathbb{E}(\zeta(\textbf{u},t))=\beta$, that could be conveniently re-parameterized as $\zeta(\textbf{u},t)=\beta+\vartheta(\textbf{u},t)$ where $\vartheta(\textbf{u},t)$ follows a zero-mean Gaussian distribution with a separable covariance structure that is temporally AR1 dependent and spatially dependent with a Matérn covariance function [Matérn, 1960], defined as:
$\text{Cov}(\vartheta(\textbf{u}_{1},t_{1}),\vartheta(\textbf{u}_{2},t_{2}))=\phi(h_{1},h_{2})=\sigma^{2}\rho(h_{1},h_{2}).$ (4)
Here, $h_{1}=\lVert\textbf{u}_{1}-\textbf{u}_{2}\rVert_{2}$, $h_{2}=|t_{1}-t_{2}|$ and $\sigma$ represents the spatio-temporal variance. In summary, the covariance function $\rho(h_{1},h_{2})$ could be defined as
$\rho(h_{1},h_{2})=\rho_{1}(h_{1})\rho_{2}(h_{2})=\left[\frac{1}{2^{\nu-1}\Gamma\left(\nu\right)}\left(\kappa\cdot h_{1}\right)^{\nu}\mathbf{K}_{\nu}\left(\kappa\cdot h_{1}\right)\right]\left[\frac{a^{h_{2}}}{1-a^{2}}\right],$ (5)
where $\nu$ is a fixed parameter, $|a|<1$ and $\kappa$ are unknown parameters, $\Gamma$ is the gamma function and $\mathbf{K}_{\nu}$ is the second-order modified Bessel function. The differentiability of the Gaussian process $\vartheta(\textbf{u},t)$ depends on the differentiability of the covariance structure; therefore, to ensure this property, we fixed a parameter $\nu\geq 3/2$.

### v2

@@ Abstract
Understanding the spread of infectious diseases such as COVID-19 is crucial for informed decision-making and resource allocation. A critical component of disease behavior is the velocity with which disease spreads, defined as the rate of change between time and space. This paper proposes a spatio-temporal modeling approach to determine the velocities of infectious disease spread. Our approach assumes that the locations and times of people infected can be considered a spatio-temporal point pattern that arises as a realization of a spatio-temporal log-Gaussian Cox point process. The intensity function of this process is estimated using a fully non-separable spatio-temporal model derived from diffusion Stochastic Partial Differential Equations (SPDE), and fast Bayesian inference is performed using integrated nested Laplace approximation (INLA). The velocity is then calculated using finite differences that approximate the derivatives of the intensity function. Finally, the directions and magnitudes of the velocities can be mapped at specific times to better examine the spread of the disease throughout the region. This method is demonstrated by analyzing COVID-19 spread in Cali, Colombia, during the 2020-2021 pandemic.
Keywords: Bayesian inference, Infectious diseases, INLA, Spatio-temporal modeling, Stochastic Partial Differential Equations
@@ 1 Introduction
Transmission of infectious diseases such as COVID-19 has the potential to collapse the healthcare system, causing health, societal, and economic problems [kaye_et_al_2021] [Ttianchen_et_al_2021]. For that reason, modeling these diseases is critical to understanding how they spread, and help decision-makers create policies to mitigate them. In this regard, a powerful tool in the context of spatio-temporal point patterns that can help understand the spread of an infectious disease is the velocity that could be computed given a direction in any location in space and time. This magnitude could be understood as the instantaneous rate of change for the point pattern at location and time $(\textbf{u},t)$ in direction v.
In the context of point patterns, velocity is a quantitative measure that helps us to understand how a spatio-temporal process evolves. Moreover, when it is possible to define the intensity function, velocity refers to how fast the intensity changes in time relative to its change in a given spatial direction. Thus, this measure captures the speed of the intensity change in that direction, and provides information about the spatio-temporal dynamics of intensity and, consequently, the underlying point pattern. In this context, velocity could be interpreted in an instantaneous sense as the chance for an event to lie in a small product neighborhood of a specific location and time.
@@
Many authors implemented different models to estimate the directional derivatives of the intensity of a point pattern. For instance, [Banerjee_2003] introduced a formal directional derivative process based on the realization of a stochastic process. Later, [terres_gelfan_2016] developed a methodology for local sensitivity analysis based on directional derivatives that also uses continuous covariates to relate them with spatial gradients. For geostatistical data, temporal and spatial derivatives were computed to estimate velocities of climate change [Schliep_et_al_2015] relating covariates to this, for example, temperature. Extending these ideas to point patterns [schliep_gelfand_2019] compute velocities for spatio-temporal point patterns, assuming that the point pattern follows a log-Gaussian Cox (LGCP) process and also by using Nearest Neighbor Gaussian process (NNGP). Additionally, other authors consider other ways to compute velocities that use higher-order derivatives, such as second-order derivatives [quick_et_al_2015].
In the context of infectious disease spreading, particularly in COVID-19 in the region of Cali, Colombia, several approaches have been used to understand, model, and predict the spread around the city. For instance, [ribeiroamaraletal23] estimates the number of infected individuals using a SIR model to predict future intensity patterns. However, this method does not provide an insightful estimation of the directions and magnitude of the disease propagation. On the other hand, [dongetal2023] uses kernel-induced feature functions to understand how fast the disease spreads around the city. However, it does not provide direction or estimation of this change. In this paper, we propose a computation of velocities for spatio-temporal point patterns using finite differences as a way to understand the magnitude and direction of disease propagation in a specific region, as it described in Section 3.3.
A key strength of our method is its flexibility: it does not assume a specific parametric form for the intensity function or a fixed covariance structure. This allows us to detect instantaneous changes that parametric models may be too rigid to capture. Nevertheless, because our nonparametric approach is data-driven, it can be sensitive to noise, bias, or irregularities in the observed data. To mitigate this, we implement the fully nonseparable covariance structure described in [lindgren2023inlaspacetime], which helps reduce data dependence and improves the consistency of velocity estimates, mainly when the number of spatio-temporal points is low, as we explain in Section 4.
For inference, we employ the integrated nested Laplace approximation (INLA) framework [rue_et_al_r_inla_2009] and inlabru [inlabru], which offers computational efficiency and simplicity for estimating the posterior distribution of the intensity function. Specifically, we use INLAspacetime [lindgren2023inlaspacetime], which extends the separable spatio-temporal covariance structure to a fully nonseparable through diffusion-based models.
@@
The structure of the remainder of the paper is as follows. In Section 2, we introduce some theory about point processes and the log-Gaussian Cox process. Then, in Section 3, we fully detail how we decompose the intensity function into three components and how inference for the LGCP is performed through its covariance structure. In this section, we also explain how velocity is defined for spatio-temporal point patterns, and how to approximate it using finite differences. In Section 4, we present a simulation study to assess the performance of our method using a point pattern with a deterministic intensity function. Then, in Section 5, we present an application using data from COVID-19 in Cali, Colombia. Finally, we present the conclusions, a discussion of the findings, and some directions for future research.
@@ 2 Preliminary concepts
In this section, we first present some theoretical concepts related to point processes necessary to introduce the methodology used to model the spread of disease and to compute velocities. Using these concepts, we can provide a detailed explanation of how disease spread can be modeled both spatially and temporally. This involves analyzing the locations of infected individuals as they change over time. By tracking these observed locations over a specific period, we can gain valuable insights into the patterns and dynamics of disease transmission. This approach allows us to understand where infections occur and how they evolve, which is crucial for effective disease control and prevention.
@@ 2.1 Point processes
Given a subset $W\subseteq\mathbb{R}^{2}$ and an interval $T\subseteq\mathbb{R}$, a spatio-temporal point pattern consists of a collection of all points $\{\zeta_{i}\}_{i=1}^{n}=\{\left(\textbf{u}_{i},t_{i}\right)\}_{i=1}^{n}\subseteq W\times T$. Under these conditions, a spatio-temporal point process is a random mechanism whose outcome or realization is a point pattern. For a spatio-temporal point process we define the intensity function $\lambda:W\times T\to[0,\infty)$ such that $\int_{A}\int_{B}\lambda(\textbf{u},t)\mathrm{d}t\mathrm{~d}\textbf{u}<\infty$ for all bounded sets $A\subseteq W$, and intervals $B\subseteq T$ that satisfies the following property
@@
One of the most important point process models is the Poisson process. This process is frequently used as a benchmark due to its simplicity, and satisfies the following conditions:
@@ 2.2 Log-Gaussian Cox processes
Log-Gaussian Cox processes (LGCPs) are typically used to model complex relationships between space and time, such as the spread of a contiguous disease [diggle_et_al_2013]. To understand this complex process, we start by discussing the Cox process. A Cox process can be viewed as a Poisson process with a random intensity function. That is, the intensity function $\Lambda(\textbf{u},t)$ meets these two postulates:
@@
    $\Lambda(\textbf{u},t)$ with $\textbf{u}\in W$ and $t\in T$ is a non-negative valued stochastic process.
@@
LGCP models were introduced as a particular case of a Cox process [mollet_et_al_1998]. As the name suggests, an LGCP is a Cox process with intensity function given by a Gaussian process, that is $\log(\Lambda(\mathbf{u},t))=\mu(\textbf{u},t)+\xi(\mathbf{u},t)$, where $\exp(\mu(\textbf{u},t))$ could be interpreted as the mean structure of $\Lambda(\textbf{u},t)$ and $\xi(\mathbf{u},t)$ is a stationary Gaussian process.
@@ 3 Methodology
To estimate velocities of an infectious disease, we propose a spatio-temporal point process model to describe how infectious diseases spread in space across time. Specifically, we use an LGCP model where the intensity is multiplicatively decomposed into temporal, spatial, and spatio-temporal stochastic variation as presented in [diggle_et_al_2005] and [ribeiroamaraletal23]. Once the model is fitted, the intensity estimate is used to compute minimal velocities through finite differences. This approach allows us to understand the direction and velocity of disease spread.
@@ 3.1 Spatio-temporal LGCP model
$\Lambda(\textbf{u},t)=\eta(\textbf{u})\cdot\mu(t)\cdot\exp\left(\xi(\textbf{u},t)\right).$ (1)
Here, the spatial and temporal components are deterministic and capture aspects of the underlying population at risk for the disease pattern. The spatial component $\eta(\textbf{u})$ represents the population density in the study region. This implies that if there were zones or locations with higher or lower cases, this variation would be absorbed in $\eta(\textbf{u})$. The temporal component $\mu(t)$ is estimated by grouping the total number of cases over time and using a spline model with different basis functions to model these data. Thus, the flexibility of the splines estimation of $\mu(t)$ captures the evolution in the number of cases over time. This emphasizes that the remaining term $\xi(\textbf{u},t)$ will capture anomalies in space and time. This term is modeled as a stationary, zero-mean log-Gaussian stochastic process. In the following sections, we explain in detail how we compute each component and their importance in the model.
@@ 3.1.1 Temporal component
The first step in our approach is to model the number of individuals per temporal unit (e.g., days, hours, etc.) infected by the disease. Accurately capturing these data is crucial for understanding the spread and making reliable predictions. We employed a spline model, a flexible statistical tool capable of fitting complex data patterns to achieve this goal.
Splines allow to handle non-linear trends in the data, providing a more precise fit than traditional linear models [wood_2017]. This flexibility is essential in epidemiological studies, where the number of infections can vary greatly due to numerous factors such as periodic changes due to day-of-week effects or time-of-year events. We therefore estimate the number of infected people in time $t$, denoted by $\mu(t)$, by fitting the following standard log-spline model
@@
where $b_{i}$ are completely known functions, commonly referred to as basis functions, and $q$ denotes the number of functions. Some examples of these basic functions are polynomials, roots, cosine, sine, and step functions. However, the election of the basic functions and their number highly depends on the data.
@@ 3.1.2 Spatial component
Different methods could be used to compute the population density $\eta(\textbf{u})$ over a study region. In this work, we use population information from WorldPop [world_pop_2020], an open-access project that combines official census information and satellite images to estimate population density at different resolutions. For regions where population information is not available, we could estimate the population density using nonparametric methods such as Gaussian kernel estimators [gonzalezandmoraga23]. For instance, the population density could be estimated as
@@
where $h>0$ is a fixed bandwidth, $\phi(\textbf{u})=(2\pi)^{-1}\exp{\{-0.5\lVert\textbf{u}\rVert^{2}\}}$, and $\textbf{u}_{i}$ are the locations of the infected individuals in the study region across time. We could also use an adaptive bandwidth $h_{i}$ for each location $\textbf{u}_{i}$ to have a better approximation.
@@ 3.1.3 Spatio-temporal component
Once the spatial and temporal components are estimated, they can be used to compute the intensity function of the spatio-temporal process. In this case, we assume that a log-Gaussian Cox process generates the observed point process. Then, the counting number of infected individuals in the region $W$ at the interval time $T$, denoted $\mathcal{N}\left(W,T\right)$, meets the following condition:
@@
where $\lambda\left(\textbf{u},t\right)$ can be decomposed as in (1) as
$\Lambda(\textbf{u},t)=\eta(\textbf{u})\cdot\mu(t)\cdot\exp\left(\xi(\textbf{u},t)\right).$ (3)
Here, the last term can be conveniently reparametrized as $\xi(\textbf{u},t)=\beta+\vartheta(\textbf{u},t)$ with $\vartheta(\textbf{u},t)$ following a zero-mean Gaussian process governed by a spatio-temporal diffusion Stochastic Partial Differential Equation (SPDE), which we introduce below. Following the approach of [lindgren2023inlaspacetime], consider the operator $L_{u}=\gamma_{u}^{2}-\Delta$, with $\gamma_{u}>0$ on a purely spatial domain $\mathcal{D}\subseteq\mathbb{R}^{d}$ on appropriated boundary conditions for compact domains. The generalized Whittle-Matérn covariance can be characterized through the precision operator $Q(\gamma_{u},\gamma_{e},\alpha)=\gamma_{e}^{2}L_{u}^{\alpha}$, where the solution $\varphi(\textbf{u},t)$ satisfies the spatial SPDE
$\gamma_{e}L_{u}^{\alpha/2}\varphi(\textbf{u})=\mathcal{W}(\textbf{u})\quad\textbf{u}\in\mathcal{D},$ (4)
where $\mathcal{W}$ is a spatial white noise process. From this equation, we can define a noise process $\text{d}\mathcal{E}_{Q}(\textbf{u},t)$ as a Gaussian noise that is uncorrelated in time but correlated in space, with precision operator $Q=Q(\gamma_{u},\gamma_{e},\alpha_{e})$ for a non-negative $\alpha_{e}$. With this noise process, these three non-negative smoothness parameters $(\alpha_{t},\alpha_{u},\alpha_{e})$ and three positive scale parameters $(\gamma_{t},\gamma_{u},\gamma_{e})$, we can define the following diffusion spatio-temporal SPDE
$\left(\gamma_{t}\frac{d}{dt}+L_{u}^{\alpha_{u}}\right)^{\alpha_{t}}v(\boldsymbol{u},t)=\mathrm{d}\mathcal{E}_{Q}(\boldsymbol{u},t),\quad(\boldsymbol{u},t)\in\mathcal{D}\times\mathbb{R}.$ (5)
The solutions $\varphi(\textbf{u},t)$ of (5) define Gaussian processes with specific structural properties. Parameter $\alpha_{e}$ governs the degree of separability in the model, where setting $\alpha_{e}=0$ leads to a fully nonseparable model. To further quantify this, [lindgren2023inlaspacetime] introduce the nonseparability parameter
$\beta_{u}=1-\frac{\alpha_{e}}{\alpha}=1-\frac{\alpha_{e}}{v_{u}+d/2},$
where $v_{u}$ represents the smoothness of the spatial process, and $\beta_{u}\in[0,1]$. The cases $\beta_{u}=0$ and $\beta_{u}=1$ correspond to fully separable and fully nonseparable models, respectively. Notably, as a special case, when $\alpha_{u}=0$, it follows that $v_{u}=\alpha_{e}-d/2$, which directly implies $\beta_{u}=0$, leading to a separable model.
Then the solution of (5) for some smoothing parameters $(\alpha_{t},\alpha_{u},\alpha_{e})$ is the residual normalized spatio-temporal component $\vartheta(\textbf{u},t)$ in (3) that captures the underlying stochastic variation in the model.